If all of the chloride in a 3.734 g sample of an unknown metal chloride is precipitated as agcl with 70.90 ml of 0.2010 m agno3 what is the percentage of chloride in the sample
1
Expert's answer
2020-10-29T05:37:58-0400

Solution:


0.2010 molar AgNO3 solution means

1000 ml of solution contain 0.2010 mole of AgNO3


1 ml of solution contain (0.2010/1000) mole AgNO3


70.90 ml of solution contain

= 70.90×(0.2010/1000) mole AgNO3

= 14.251×10^-3 mole AgNO3


AgNO3 react with Cl according to the following equation:

AgNO3 + Cl ————> AgCl + NO3

1mole of AgNO3 is required to fully reacts with 1 mole of Cl

So,

If 14.251×10-3 mole AgNO3 react then there must be 14.251×10-3 mole Clin the given 3.734 g of unknown metal chloride


molar mass of Cl is = 35.5 g

14.251×10–3 mole Cl has mass

=(35.5×14.251×10^-3) g

= 0.506 g



So, percent of chloride in 3.734 g of metal chloride

= (0.506 g/3.734 g)×100 %

= 13.548 %


Hence, 13.548 % of chloride present in 3.734 g of unknown metal chloride